OpenStudy (thecatman):

Which set of line segments with the given lengths can be formed into a right triangle? 24, 30, 35 12, 18, 30 18, 24, 30 18, 24, 35

OpenStudy (thecatman):

i think c or d

OpenStudy (mathstudent55):

See if each option works in the Pythagorean theorem equation. If it does, it is a right triangle. \(a^2 + b^2 = c^2\) Let the largest side be c, and use the other two sides as a and b.

OpenStudy (thecatman):

ok thank you

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

I suggest the other way to find out :) as you know, 3,4,5 is the right triangle for sure, right? let say a =3 b=4 c =5 therefore, a= \(\dfrac{3}{4}b\) =\(\dfrac{3}{5}c\) so that, I just check the equation above for example: 18 =\(\dfrac{3}{4}24 \) \(\surd\) and then 18 =\(\dfrac{3}{5}30\) \(\surd\) --> 18,24,30 is the answer